Home‎ > ‎Crew 133‎ > ‎

117-ScienceReport

Science Report
01/17/2014
Paula Crock

EVA helmet temps:  Data were collected during a total of 11 EVAs.  The devices seemed to work properly.  Troubleshooting over the past two weeks has refined the data collection method and helped to establish a secure, non-marring method for attaching the devices to the EVA backpacks. Detailed instructions including clear IRB protocol instructions have been written and will be provided for crew members through the rest of the season who may wish to participate in this study. Equipment, paperwork and instructions will be left in the lab.  Mailing packaging and prepaid postage will be supplied at the end of the season to mail everything back.

Radio Dead Zone map:  Data was collected on walking EVAs west, east and north of the Hab and on ATV EVAs north of the Hab.  Data consists of lat/lon coordinates and success/failure of radio communication with Hab and, if success, whether the communication was clear.  Data were collected on 22 data points.  This data will be incorporated into the model-generated radio "dead zone" map that is still in development.  The ground truth data points will be used to refine/validate the model. The ground truth part of this study could be extended to other crews through the rest of the season with minimal instruction that I would be happy to provide.